# Arveson’s hyperrigidity conjecture is false

### v1

@@ 1. Introduction
Let $G$ be a separable subset of bounded operators on Hilbert space generating a norm-closed self-adjoint unital subspace $\mathcal{S}$, which in turn generates a C*-algebra $\mathcal{B}$. Arveson’s hyperrigidity conjecture [8, Conjecture 4.3] states that if the restriction to $\mathcal{S}$ of all irreducible ***-representations of $\mathcal{B}$ have unique completely positive extensions to $\mathcal{B}$, then $G$ is hyperrigid in the sense that for every faithful ***-representation on Hilbert space $\mathcal{B}\subseteq\mathbb{B}(\mathcal{H})$, and a sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ of unital completely positive maps on $\mathbb{B}(\mathcal{H})$ we have,
@@
There is an operator $T$ on Hilbert space such that the the unital operator algebra $\mathcal{A}$ generated by $T$ is not hyperrigid in $\mathcal{B}=C^{*}(\mathcal{A})$, while the restriction to $\mathcal{A}$ of all irreducible ***-representations of $\mathcal{B}$ have unique completely positive extensions to $\mathcal{B}$. In fact, $T$ is a rank-one perturbation of a unitary, so that with $G=\mathcal{A}$ we get a counterexample to Arveson’s hyperrigidity conjecture with $\mathcal{B}$ of type I.
Hyperrigidity was first defined by Arveson in [8] as a non-commutative analogue of a rigidity of approximations in classical approximation theory [30, 34, 10] (see [18, Definition 5.2]). In [8, Theorem 2.1] Arveson showed that a unital operator algebra $\mathcal{A}$ is hyperrigid in $\mathcal{B}$ if and only if every ***-representation $\pi$ of the generated C*-algebra $\mathcal{B}$ has a unique completely positive extension when restricted to $\mathcal{A}$. Irreducible ***-representations $\pi$ for which $\pi|_{\mathcal{A}}$ has this property are called boundary representations, and together they form the non-commutative Choquet boundary. The non-commutative Choquet boundary is a generalization of classical Choquet boundary of a uniform algebra, defined as the set of points with unique representing measures (see for instance [24]).
@@
Theorem 1.1 illustrates that in order to obtain a non-commutative generalization of Šaškin’s theorem, hyperrigidity will have to be replaced with a different kind of rigidity of approximations. Our counterexample was discovered by attempting to extend results of the second-named author with Salomon [21] and describe all boundary representations of tensor algebras in the sense of Muhly and Solel [31]. These results will appear in a separate forthcoming paper.
@@ 2. Background
We discuss some of the background on non-commutative boundary theory. We refer the reader to [33, 9] for some of the necessary prerequisites needed for our paper. A unital operator algebra $\mathcal{A}$ is a norm-closed unital subalgebra of bounded operators on Hilbert space, and we denote by $\mathcal{B}=C^{*}(\mathcal{A})$ the C*-algebra generated by $\mathcal{A}$. A representation $\rho:\mathcal{A}\rightarrow\mathbb{B}(\mathcal{H})$ of $\mathcal{A}$ is a unital completely contractive homomorphism. We say a representation $\rho$ has the unique extension property if there is a unique completely positive extension of $\rho$ to $\mathcal{B}$, and this extension is multiplicative. An irreducible ***-representation $\pi$ of $\mathcal{B}$ is called a boundary representation if $\pi|_{\mathcal{A}}$ has the unique extension property.
For the purpose of finding a counterexample to Arveson’s hyperrigidity conjecture, we may restrict our attention to unital operator algebras instead of operator systems since any representation of a unital operator algebra $\mathcal{A}$ extends uniquely to a unital completely positive map on the operator system $\mathcal{S}=\overline{\mathcal{A}+\mathcal{A}^{*}}$. Hence, a counterexample of a separable unital operator algebra would provides us with a counterexample of a separable operator system generating the same C*-algebra. Moreover, we also have by [8, Theorem 2.1] that a separable unital operator algebra $\mathcal{A}$ is hyperrigid in its generated C*-algebra $\mathcal{B}$ if and only if for every ***-representation $\pi$ of $\mathcal{B}$, the restriction $\pi|_{\mathcal{A}}$ has the unique extension property.
@@ 3. A Type I counterexample
Lemma 3.1.
For any $\xi\in\mathcal{H}$ we have that $\langle T^{n}\xi,\xi\rangle\underset{n\rightarrow\infty}{\longrightarrow}0$.
Proof.
We may write any $\xi\in\mathcal{H}$ as $\xi=c\cdot\eta+\zeta$ for a scalar $c\in\mathbb{C}$ and $\zeta=\sum_{m\in\mathbb{Z}}a_{m}\cdot e_{m}\in\ell^{2}(\mathbb{Z})$ for scalars $a_{m}\in\mathbb{C}$ with $\sum_{m\in\mathbb{Z}}|a_{m}|^{2}<\infty$. Since $T\eta=e_{0}$ and $Te_{m}=e_{m+1}$ for all $m\in\mathbb{Z}$, we see that for $n\geq 1$,
$\langle T^{n}\xi,\xi\rangle=|c|^{2}\cdot\langle T^{n}\eta,\eta\rangle+c\cdot\langle T^{n}\eta,\zeta\rangle+\overline{c}\cdot\langle U^{n}\zeta,\eta\rangle+\langle U^{n}\zeta,\zeta\rangle$
$=c\cdot a_{n-1}+\langle U^{n}\zeta,\zeta\rangle\underset{n\rightarrow\infty}{\longrightarrow}0,$
because $a_{n-1}\rightarrow 0$, and $U^{n}\rightarrow 0$ in the weak operator topology. ∎
@@
Let $q:\mathcal{B}\rightarrow C(\mathbb{T})$ be the natural quotient map sending $T$ to the identity function $z\mapsto z$ on $\mathbb{T}$. Thus, by [4, Theorem 1.3.4] and the paragraph preceding it, up to unitary equivalence the irreducible ***-representations of $\mathcal{B}$ are comprised of the identity representation $\iota:\mathcal{B}\rightarrow\mathbb{B}(\mathcal{H})$ (see [4, Page 20, Corollary 2]), and evaluations $\tau_{z}:=ev_{z}\circ q:\mathcal{B}\rightarrow\mathbb{B}(\mathbb{C})$ by unimodular scalars $z\in\mathbb{T}$.
@@
$\mathcal{A}$ is generated by $T$ and we saw already that $\mathcal{B}=C^{*}(T)$ is type I. Thus, we need to show that $\mathcal{A}$ is not hyperrigid in $\mathcal{B}$ while the restriction of $\mathcal{A}$ of all irreducible ***-representations of $\mathcal{B}$ have the unique extension property.
First, we show that $\mathcal{A}$ is not hyperrigid in $\mathcal{B}$. Due to the equivalence between maximality and the unique extesnion property, it suffices to exhibit a ***-representation $\sigma$ of $\mathcal{B}$ such that $\sigma|_{\mathcal{A}}$ admits a non-trivial dilation. Let $\tau:C(\mathbb{T})\rightarrow\mathbb{B}(\ell^{2}(\mathbb{Z}))$ be the ***-representation of the bilateral shift given by $z\mapsto U$ which lifts to a ***-representation $\sigma:=\tau\circ q$ of $\mathcal{B}$. Then, by construction we have for all $A\in\mathcal{A}$ that
@@
Second, we show that all irreducible ***-representations of $\mathcal{B}$ are boundary representations for $\mathcal{A}$. It is clear that the quotient map $q:\mathcal{B}\rightarrow C(\mathbb{T})$ is not isometric on $\mathcal{A}$ because the norm of $T$ becomes strictly smaller. Hence, by Arveson’s boundary theorem (see for instance [16]) we have that the identity representation $\iota$ is a boundary representation.
Now, let $z\in\mathbb{T}$ be a unimodular scalar. We will show that $\tau_{z}:\mathcal{B}\rightarrow\mathbb{B}(\mathbb{C})$ is a boundary representation by showing that $\tau_{z}|_{\mathcal{A}}$ is maximal. By the Dritschel–McCullough dilation theorem [23] (see also [9]) we know that $\tau_{z}|_{\mathcal{A}}$ admits a dilation to a maximal representation $\rho:\mathcal{A}\rightarrow\mathbb{B}(\mathcal{L})$ with $\xi\in\mathcal{L}$ a unit vector which is the image of $1$ under the dilating isometry $V:\mathbb{C}\rightarrow\mathcal{L}$. Let $\pi$ be the extension of $\rho$ to a ***-representation, guaranteed by the unique extension property for $\rho$. We assume without loss of generality that $\mathcal{L}$ is the smallest reducing subspace containing $\rho(\mathcal{A})\xi$. Then, by the paragraph preceding [4, Theorem 1.3.4] we have that $\pi=\pi_{\mathbb{K}}\oplus\pi_{C(\mathbb{T})}$ where $\pi_{\mathbb{K}}$ is a multiple of the identity representation of $\mathcal{B}$ (see [4, Page 20, Corollary 1]), and $\pi_{C(\mathbb{T})}$ is a representation that annihilates the ideal of compact operators $\mathbb{K}(\mathcal{H})$. Let $\mathcal{L}_{\mathbb{K}}\oplus\mathcal{L}_{\mathbb{T}}$ be the orthogonal decomposition of $\mathcal{L}$ with respect to the decomposition $\pi=\pi_{\mathbb{K}}\oplus\pi_{C(\mathbb{T})}$, and let $Q$ and $R$ be the orthogonal projections onto $\mathcal{L}_{\mathbb{K}}$ and $\mathcal{L}_{\mathbb{T}}$ respectively.
On the one hand, since $\rho$ dilates $\tau_{z}$ on $\mathcal{A}$ we have that $|\langle\rho(T^{n})\xi,\xi\rangle|=|\tau_{z}(T^{n})|=1$ for all $n\in\mathbb{N}$. On the other hand if $Q\xi\neq 0$, since $\rho(T^{n})$ on $\mathcal{L}_{\mathbb{K}}=Q\mathcal{L}$ is unitarily equivalent to a multiple of $T^{n}$ on $\mathcal{H}$ we get by Lemma 3.1 that $\langle\rho(T^{n})Q\xi,Q\xi\rangle\rightarrow 0$. Hence,
@@
where the last strict inequality follows by Pythagoras’s theorem, and the second to last inequality follows from Cauchy-Schwarz inequality together with the fact that $R\rho(T^{n})R=\pi_{C(\mathbb{T})}(T^{n})$ is a contraction. Thus, since $\mathcal{L}$ is the smallest reducing subspace containing $\rho(\mathcal{A})\xi$, we must have that $Q\xi=0$ and $\mathcal{L}=\mathcal{L}_{\mathbb{T}}$, so that $\pi=\pi_{C(\mathbb{T})}$ annihilates $\mathbb{K}(\mathcal{H})$. But now, since $\pi(T)=\rho(T)$ is a unitary dilation of the unitary $\tau_{z}(T)$, we must have that $\tau_{z}|_{\mathcal{A}}$ is a direct summand of $\rho$. Thus, $\tau_{z}|_{\mathcal{A}}=\rho$ is maximal, and $\tau_{z}$ is a boundary representation for all $z\in\mathbb{T}$. ∎
@@
Acknowledgments. The authors are grateful to Kevin Aguyar Brix, Raphaël Clouâtre, Ken Davidson, Travis Russell, Guy Salomon and Orr Shalit for remarks on preliminary versions of this paper. We also thank Lucas Hall for organizing the Dilation Theory learning seminar which took place at the Technion in the Fall semester of 2023-2024.

### v2

@@ 1. Introduction
Let $G$ be a separable subset of bounded operators on Hilbert space generating a norm-closed self-adjoint unital subspace $\mathcal{S}$, which in turn generates a C*-algebra $\mathcal{B}$. Arveson’s hyperrigidity conjecture [8, Conjecture 4.3] states that if the restriction to $\mathcal{S}$ of all irreducible *-representations of $\mathcal{B}$ have unique completely positive extensions to $\mathcal{B}$, then $G$ is hyperrigid in the sense that for every faithful *-representation on Hilbert space $\mathcal{B}\subseteq\mathbb{B}(\mathcal{H})$, and a sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ of unital completely positive maps on $\mathbb{B}(\mathcal{H})$ we have,
@@
There is an operator $T$ on Hilbert space such that the the unital operator algebra $\mathcal{A}$ generated by $T$ is not hyperrigid in $\mathcal{B}=C^{*}(\mathcal{A})$, while the restriction to $\mathcal{A}$ of all irreducible *-representations of $\mathcal{B}$ have unique completely positive extensions to $\mathcal{B}$. In fact, $T$ is a rank-one perturbation of a unitary, so that with $G=\mathcal{A}$ we get a counterexample to Arveson’s hyperrigidity conjecture with $\mathcal{B}$ of type I.
Hyperrigidity was first defined by Arveson in [8] as a non-commutative analogue of a rigidity of approximations in classical approximation theory [30, 34, 10] (see [18, Definition 5.2]). In [8, Theorem 2.1] Arveson showed that a unital operator algebra $\mathcal{A}$ is hyperrigid in $\mathcal{B}$ if and only if every *-representation $\pi$ of the generated C*-algebra $\mathcal{B}$ has a unique completely positive extension when restricted to $\mathcal{A}$. Irreducible *-representations $\pi$ for which $\pi|_{\mathcal{A}}$ has this property are called boundary representations, and together they form the non-commutative Choquet boundary. The non-commutative Choquet boundary is a generalization of classical Choquet boundary of a uniform algebra, defined as the set of points with unique representing measures (see for instance [24]).
@@
Theorem 1.1 illustrates that in order to obtain a non-commutative generalization of Šaškin’s theorem, hyperrigidity will have to be replaced with a different kind of rigidity of approximations. We discovered our counterexample by attempting to extend results of the second-named author with Salomon [21] and describe all boundary representations of tensor algebras in the sense of Muhly and Solel [31]. These results will appear in a separate forthcoming paper.
@@ 2. Background
We discuss some of the background on non-commutative boundary theory. We refer the reader to [33, 9] for some of the necessary prerequisites needed for our paper. A unital operator algebra $\mathcal{A}$ is a norm-closed unital subalgebra of bounded operators on Hilbert space, and we denote by $\mathcal{B}=C^{*}(\mathcal{A})$ the C*-algebra generated by $\mathcal{A}$. A representation $\rho:\mathcal{A}\rightarrow\mathbb{B}(\mathcal{H})$ of $\mathcal{A}$ is a unital completely contractive homomorphism. We say a representation $\rho$ has the unique extension property if there is a unique completely positive extension of $\rho$ to $\mathcal{B}$, and this extension is multiplicative. An irreducible *-representation $\pi$ of $\mathcal{B}$ is called a boundary representation if $\pi|_{\mathcal{A}}$ has the unique extension property.
For the purpose of finding a counterexample to Arveson’s hyperrigidity conjecture, we may restrict our attention to unital operator algebras instead of operator systems since any representation of a unital operator algebra $\mathcal{A}$ extends uniquely to a unital completely positive map on the operator system $\mathcal{S}=\overline{\mathcal{A}+\mathcal{A}^{*}}$. Hence, a counterexample of a separable unital operator algebra would provides us with a counterexample of a separable operator system generating the same C*-algebra. Moreover, we also have by [8, Theorem 2.1] that a separable unital operator algebra $\mathcal{A}$ is hyperrigid in its generated C*-algebra $\mathcal{B}$ if and only if for every *-representation $\pi$ of $\mathcal{B}$, the restriction $\pi|_{\mathcal{A}}$ has the unique extension property.
@@ 3. A Type I counterexample
Lemma 3.1.
For any set $I$ and two vectors $(\xi_{i}),(\eta_{i})\in\oplus_{i\in I}\mathcal{H}$ we have,
$\sum_{i\in I}\langle T^{n}\xi_{i},\eta_{i}\rangle=\langle(\oplus_{i\in I}T^{n})(\xi_{i}),(\eta_{i})\rangle\underset{n\rightarrow\infty}{\longrightarrow}0.$
Proof.
To prove the lemma, it is equivalent to show that $\{T^{n}\}_{n\in\mathbb{N}}$ converges to 00 in the $\sigma$-weak topology. We abuse notation and consider $U$ as an operator on $\mathcal{H}$ by extending it to be 00 on $\mathbb{C}\eta$. Then, we may write $T^{n+1}=U^{n}T$, so that $\|T^{n}\|=\sqrt{2}$ for all $n\in\mathbb{N}$ and the sequence $\{T^{n}\}_{n\in\mathbb{N}}$ is bounded. Since the $\sigma$-weak and weak operator topologies coincide on bounded sets, it will suffice to show that $\{T^{n}\}_{n\in\mathbb{N}}$ converges to 00 in the weak operator topology. Let $\xi,\eta\in\mathcal{H}$, and write
$\langle T^{n+1}\xi,\eta\rangle=\langle U^{n}T\xi,\eta\rangle=\langle U^{n}T\xi,(1-P)\eta\rangle.$
Since $\{U^{n}\}_{n\in\mathbb{N}}$ on $\ell^{2}(\mathbb{Z})$ converges to 00 in the weak operator topology, by taking $\xi^{\prime}:=T\xi$ and $\eta^{\prime}:=(1-P)\eta$ we get that $\langle U^{n}\xi^{\prime},\eta^{\prime}\rangle\rightarrow 0$. Hence, $\{T^{n}\}_{n\in\mathbb{N}}$ converges to 00 in the $\sigma$-weak topology. ∎
@@
Let $q:\mathcal{B}\rightarrow C(\mathbb{T})$ be the natural quotient map sending $T$ to the identity function $z\mapsto z$ on $\mathbb{T}$. According to [4, Theorem 1.3.4] and the paragraph preceding it, the spectrum of $\mathcal{B}$ is made up of the identity representation $\iota:\mathcal{B}\rightarrow\mathbb{B}(\mathcal{H})$ (see [4, Page 20, Corollary 2]), and evaluations $\tau_{z}:=ev_{z}\circ q:\mathcal{B}\rightarrow\mathbb{B}(\mathbb{C})$ by unimodular scalars $z\in\mathbb{T}$.
@@
$\mathcal{A}$ is generated by $T$ and we saw already that $\mathcal{B}=C^{*}(T)$ is type I. Thus, we need to show that $\mathcal{A}$ is not hyperrigid in $\mathcal{B}$ while the restriction of $\mathcal{A}$ of all irreducible *-representations of $\mathcal{B}$ have the unique extension property.
First, we show that $\mathcal{A}$ is not hyperrigid in $\mathcal{B}$. Due to the equivalence between maximality and the unique extesnion property, it suffices to exhibit a *-representation $\sigma$ of $\mathcal{B}$ such that $\sigma|_{\mathcal{A}}$ admits a non-trivial dilation. Let $\tau:C(\mathbb{T})\rightarrow\mathbb{B}(\ell^{2}(\mathbb{Z}))$ be the *-representation of the bilateral shift given by $z\mapsto U$ which lifts to a *-representation $\sigma:=\tau\circ q$ of $\mathcal{B}$. Then, by construction we have for all $A\in\mathcal{A}$ that
@@
Second, we show that all irreducible *-representations of $\mathcal{B}$ are boundary representations for $\mathcal{A}$. It is clear that the quotient map $q:\mathcal{B}\rightarrow C(\mathbb{T})$ is not isometric on $\mathcal{A}$ because the norm of $T$ becomes strictly smaller. Hence, by Arveson’s boundary theorem (see for instance [16]) we have that the identity representation $\iota$ is a boundary representation.
Now, let $z\in\mathbb{T}$ be a unimodular scalar. We will show that $\tau_{z}:\mathcal{B}\rightarrow\mathbb{B}(\mathbb{C})$ is a boundary representation by showing that $\tau_{z}|_{\mathcal{A}}$ is maximal. By the Dritschel–McCullough dilation theorem [23] (see also [9]) we know that $\tau_{z}|_{\mathcal{A}}$ admits a dilation to a maximal representation $\rho:\mathcal{A}\rightarrow\mathbb{B}(\mathcal{L})$ with $\xi\in\mathcal{L}$ a unit vector which is the image of $1$ under the dilating isometry $V:\mathbb{C}\rightarrow\mathcal{L}$. Let $\pi$ be the extension of $\rho$ to a *-representation, guaranteed by the unique extension property for $\rho$. We assume without loss of generality that $\mathcal{L}$ is the smallest reducing subspace containing $\rho(\mathcal{A})\xi$. Then, by the paragraph preceding [4, Theorem 1.3.4] we have that $\pi=\pi_{\mathbb{K}}\oplus\pi_{C(\mathbb{T})}$ where $\pi_{\mathbb{K}}$ is a multiple of the identity representation of $\mathcal{B}$ (see [4, Page 20, Corollary 1]), and $\pi_{C(\mathbb{T})}$ is a representation that annihilates the ideal of compact operators $\mathbb{K}(\mathcal{H})$. Let $\mathcal{L}_{\mathbb{K}}\oplus\mathcal{L}_{\mathbb{T}}$ be the orthogonal decomposition of $\mathcal{L}$ with respect to the decomposition $\pi=\pi_{\mathbb{K}}\oplus\pi_{C(\mathbb{T})}$, and let $Q$ and $R$ be the orthogonal projections onto $\mathcal{L}_{\mathbb{K}}$ and $\mathcal{L}_{\mathbb{T}}$ respectively.
On the one hand, since $\rho$ dilates $\tau_{z}$ on $\mathcal{A}$ we have that $|\langle\rho(T^{n})\xi,\xi\rangle|=|\tau_{z}(T^{n})|=1$ for all $n\in\mathbb{N}$. On the other hand if $Q\xi\neq 0$, since the sequence $\{\rho(T^{n})\}_{n\in\mathbb{N}}$ on $\mathcal{L}_{\mathbb{K}}=Q\mathcal{L}$ is unitarily equivalent to a multiple of $\{T^{n}\}_{n\in\mathbb{N}}$ on $\mathcal{H}$, by Lemma 3.1 we get $\langle\rho(T^{n})Q\xi,Q\xi\rangle\rightarrow 0$. Hence,
@@
where the last strict inequality follows by the Pythagorean theorem, and the second to last inequality follows from Cauchy-Schwarz inequality together with the fact that $R\rho(T^{n})R=\pi_{C(\mathbb{T})}(T^{n})$ is a contraction. Thus, since $\mathcal{L}$ is the smallest reducing subspace containing $\rho(\mathcal{A})\xi$, we must have that $Q\xi=0$ and $\mathcal{L}=\mathcal{L}_{\mathbb{T}}$, so that $\pi=\pi_{C(\mathbb{T})}$ annihilates $\mathbb{K}(\mathcal{H})$. But now, since $\pi(T)=\rho(T)$ is a unitary dilation of the unitary $\tau_{z}(T)$, we must have that $\tau_{z}|_{\mathcal{A}}$ is a direct summand of $\rho$. Thus, $\tau_{z}|_{\mathcal{A}}=\rho$ is maximal, and $\tau_{z}$ is a boundary representation for all $z\in\mathbb{T}$. ∎
@@
Acknowledgments. The authors are grateful to Kevin Aguyar Brix, Raphaël Clouâtre, Lucas Hall, Michael Hartz, Ken Davidson, Travis Russell, Guy Salomon and Orr Shalit for comments, remarks and suggestions on preliminary versions of this paper. We especially thank Lucas Hall for organizing the Dilation Theory learning seminar which took place at the Technion in the Fall semester of 2023-2024.